OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a15 = -53, a16 = -5

OpenStudy (anonymous):

an = -725 + 48(n - 1)

OpenStudy (campbell_st):

well you know the general term is \[a_{n} = a + (n - 1) \times d\] a = 1st term, n = number of terms and d = common difference well you know the difference between the 15th and 16th term is 48 so the common difference d = 48 substitute it to find a. I've used the 15th term \[-53 = a + (15 -1) \times 48\] so -53 = a + 672 then a = -725 so the general term is \[a_{n} = -725 + (n - 1) \times 48\] you'll need to distribute the 48 and then collect like terms to get the final version of the general term. hope this helps

OpenStudy (anonymous):

thnx